# library.bib

### v1

@@ Abstract
Collectives are often able to process information in a distributed fashion, surpassing each individual member’s processing capacity. In fission-fusion dynamics, where group members come together and split from others often, sharing complementary information about uniquely known foraging areas could allow a group to track a heterogenous foraging environment better than any group member on its own. We analyse the partial overlaps between individual core ranges, which we assume represent the knowledge of an individual during a given season. We identify sets of individuals whose overlap shows a balance between redundantly and uniquely known portions and we use simplicial complexes to represent these higher-order interactions. The structure of the simplicial complexes shows holes in various dimensions, revealing complementarity in the foraging information that is being shared. We propose that the complex spatial networks arising from fission-fusion dynamics allow for adaptive, collective processing of foraging information in dynamic environments.
@@ Introduction
A driver of the structure and dynamics of networked systems is their adaptive capacity to collectively process information in ways that none of the components would be able to alone [mitchell2006complex]. A prevalent feature of this information processing is its distributed nature: different components complement one another in the information they contain, which they aggregate via the network of interactions. This complementarity in information aggregation is a hallmark of collectively intelligent systems [bettencourt2009rules, garg2025causal].
@@
In social animals, the sharing of information about resource availability between group members has long been considered one of the benefits of living in groups [clark1984foraging, galef2001social, valone2002public]. Information transfer between individuals allows the group as a whole to gain better knowledge of the location of available food sources than any individual would gain on its own, especially in heterogenous, dynamic environments [clark1984foraging, kummer2006primate]. This information transfer need not be through the use of active signals, and may simply require that naïve individuals follow knowledgeable ones to find food sources that the latter know uniquely (e.g. [sonerud2001ignorant, petit2010decision, king2011next]). In species with flexible grouping patterns and wide foraging areas, knowledge could be distributed among all members of the group such that they may act as naïve or knowledgeable with respect to different food sources [sonerud2001ignorant, palacios2019uncovering]. These dynamics may allow the benefits of being in a group to outweigh the costs of competition caused by foraging with others [clark1984foraging].
@@
Fission-fusion dynamics are a property of many animal groups [aureli2008fission, ramos2018fission] that occupy a common area but are often divided into subgroups that constantly come together and split from each other. In species with a high degree of these dynamics, the entire group is never found together, and subgroups of varying sizes spread over large areas, travelling independently of other subgroups [ramos2011no], but also coming together often [pinacho2017influence]. In terms of group-decision making, fission-fusion dynamics can be seen as the result of a lack of consensus between individuals who may have different priorities on their activities [sueur2011collective]. However, in heterogenous, dynamic environments, this lack of consensus may have advantages in terms of distributed foraging and information gathering: individuals may forage in different regions of the group home range, while in subgroups with different combinations of individuals [ramos2018quantifying]. Thus, a subgroup will often contain individuals who have been foraging in (and perhaps have knowledge about) different regions of their home range.
@@
Spider monkeys (Ateles spp) are frugivorous primates living in tropical forests, where fruit can be found in ephemeral patches that vary in time and space depending on the tree species [difiore2008diets]. They have a high degree of fission-fusion dynamics [ramos2014unraveling] in terms of the variation in subgroup size [pinacho2017influence] and composition [ramos2018quantifying], and they occupy partially overlapping home ranges [smith2016seasonal]. In previous work [palacios2019uncovering], we showed that spider monkeys share information about newly discovered fruiting trees. Focusing our observations on relevant trees during their whole fruiting period, we showed that individuals that were naïve about the presence of fruit in the focal tree tended to arrive with other group members that already knew about it. Using a random arrival null model, we also showed that the group as a whole finds out about the fruiting tree in fewer visits than if each individual had to find it on its own [palacios2019uncovering].
@@
Partly inspired by those results, Falcón-Cortés et al. [falcon2019collective] used an agent-based model to explore the effects of information sharing and memory of visited patches on the collective estimation of food patch quality. When individual agents remember previously visited patches, move toward them depending on their quality, and copy others’ visiting patterns when not finding food on their own, the group as a whole is able to compute the relative rank of food patches in the environment, concentrating its activity in proportion to the size of patches and focusing foraging efforts around the largest patches [falcon2019collective]. These studies show that sharing complementary information about heterogenous environments can effectively lead to an improvement in the group’s overall knowledge, and thus could be considered a form of collective information processing, sensu [moussaid2009collective].
@@
Here we assume that information sharing takes place between individuals, in the context of a fluid grouping pattern where knowledgeable and naïve individuals coincide temporarily in the same subgroups and locations, being able to share information with others about areas they uniquely know. We also assume that an individual’s core range represents, for a given season, the area where it knows the location of available fruiting trees relatively well. This assumption is supported by the species’ wide daily travel distance (between 500 and 4500m with means between 1750 and 3311m in different study sites: [wallace2008factors]) when an individual visits around 10 different fruiting trees in a single day [suarez2006diet]. Individual core ranges vary across seasons as would be predicted based on fruit availability, with dry seasons showing larger and less overlapping core ranges than wet seasons [smith2016seasonal].
@@
The way in which individual core ranges overlap shows that, normally, more than two individuals (but not the whole group) coincide in most portions of a group’s range [smith2016seasonal], such that range overlaps can be thought of as higher-order interactions between two or more individuals. Therefore, typical network approaches to describe the full space-sharing structure of the group are inadequate due to their inability to capture multibody (non-dyadic) interactions between individuals [Silk2022, silk2023conceptual]. Simplicial sets are increasingly applied to model such scenarios where non-dyadic interactions are key features of network structure [Iacopini2024, Iacopini2024Temporal, hasenjager2024social, Lin2024]. Simplicial sets are generalisations of ordinary networks which allow for arbitrarily many nodes or individuals to be connected by a single higher-order edge called a simplex (Figure 1). For spatial problems, a particularly relevant class of simplicial structure is the simplicial complex [Silk2022] which imposes the additional assumption of downward closure; meaning that the simplicial set must contain all possible nested lower-order simplices of any simplex [silk2023conceptual, Serrano2020, Otter2017]. For example, a simplicial complex with a triadic connection between three individuals will also contain the three dyadic connections between these individuals. This assumption of downward closure provides a suite of additional computational tools for studying higher-order structure [Otter2017].
@@
Simplicial complex structure is often described through the computation of Betti numbers. There is a Betti number for each dimension, denoted $\beta_{i}$, which gives the number of holes of a particular simplex in the $i$-th dimension. For instance, $\beta_{0}$ is the number of connected components, $\beta_{1}$ the number of loops and $\beta_{2}$ the number of voids [Otter2017]. For higher dimensions, we cannot visualise the holes, so the Betti numbers are especially useful to describe higher dimensional topological structure. In the context of a higher-order network structure of spatial overlaps, the topological holes described by the Betti numbers may characterise complementarity in the information possessed by different subgroups (represented here by simplices). In addition, simplicial centrality measures [Serrano2020] allow the identification of simplices that are well connected to others in the simplicial complex.
@@
We hypothesise that, by pooling complementary information about the location of feeding trees, the group as a whole acquires a more complete knowledge of a complex, dynamic environment during a given season, than any individual on its own. In particular, given our knowledge of spider monkey socio-spatial ecology, we expect to find evidence of complementarity in the structure of the simplicial complex representing the overlaps of the individual core ranges. Our strategy, outlined in Figure 1, involves first calculating individual core ranges and their relative intersections. We assume that a set of individuals will share complementary information as a result of a balance between their redundant and unique knowledge (or between the intersection and the union of their areas). The intersection is where they can coincide to share information, while the union includes areas known uniquely by some, which they can share with the rest.
@@
We use varying thresholds of this ratio, connecting sets of individuals with decreasingly redundant overlaps into simplicial complexes. Then, we explore the structure of these complexes, searching for holes in various dimensions (given by the Betti numbers) and their persistence for different thresholds of redundancy. We interpret these holes and their persistence as evidence of complementarity in the foraging information sharing by the whole network. We also compare the simplicial complexes between the dry and wet seasons, as the uncertainty about the foraging environment may be greater in the former. We expected to find evidence for more complementarity in the information being shared during the dry seasons. Finally, we analyse the centrality patterns in the structure of the simplicial complex, to identify those sets of individuals that are key in terms of sharing foraging information with others by participating in more simplices.
@@
Figure 1: Strategy to analyse spatial overlaps in individual core ranges as higher-order interactions in a simplicial complex. In the upper left, four hypothetical areas from individuals $A$, $B$, $C$ and $D$ overlap to different extents. The balance of a spatial interaction between a set of any size, $n$, is defined by the intersection/union ratio, which scales with $n$ as $1/(n+1)$, a relationship that we capture with the filtration index, $\alpha$ (see Methods for details). Increasing values of $\alpha$ imply the consideration of decreasingly redundant spatial interactions to connect sets of individual core ranges. In the hypothetical example, the areas $A$, $B$ and $C$ have a more redundant spatial interaction than the areas $A$, $B$ and $D$, which have a smaller redundant component compared to the uniquely known portions of their union. Thus, the simplex $[A,B,C]$ is created at lower $\alpha$ values than the simplex $[A,B,D]$. In the sequence of simplices in the middle, we show different configurations of the resulting simplicial complex, adding links depending on the redundancy of the spatial interaction. Going from left to right, assuming no interactions would represent these four areas as nodes in an unconnected graph, or a 0-simplex. As we incorporate spatial interactions according to the balance between redundant and unique knowledge, we introduce links that join the initially disconnected nodes. First the link between $A$ and $B$, who intersect the most, then $B$ and $C$, later $A$ and $C$. As these simplices appear and close a circuit, they leave a hole of dimension 1 which is filled by a bi-dimensional triangle when our filtering includes the interaction between the three of them. Finally, when incorporating the smaller and less redundant spatial interaction between $D$ and the rest, we obtain a 3-simplex which contains all simplices within it, including the three-dimensional space inside the tetrahedron, which would represent the four-way interaction between all nodes. Below each simplex we note the simplices it contains (in brackets), and the highest order of the resulting simplex in the bottom. The barcode on top represents the persistence of the four connected components as we incorporate less redundant and more unique interactions (red bars) and the persistence of the 1-dimensional hole (blue bar) which appears when connecting nodes $A$ and $C$ and disappears when we connect the three nodes $A$, $B$ and $C$ by a three-way interaction. The number of red and blue bars at a given point in the filtration represents the Betti numbers $\beta_{0}$ and $\beta_{1}$, respectively.
@@ Results
Figure 2: Maps of overlaid individual core ranges, for different years and seasons (top row: dry seasons; bottom row: wet seasons). Each polygon filled with a different colour shade corresponds to an individual (scale bar in the lower right). Below each map is the number of individual core ranges overlaid, as well as the area comprised by their union, in hectares.
Figure 3 shows the density distribution of individual core ranges for the different years and seasons. Individual ranges differed significantly between years and seasons (ANOVA interaction between years and seasons, $F_{1,135}$=1.94, $p$=0.09; year: $p$<0.001; season: $p$ <0.001). During a given year, individual spider monkeys occupied larger core ranges during the dry season than during the wet season. This is particularly evident in 2014, 2015 and 2016.
@@
We developed a theoretical prediction of the extent to which the core ranges of a set of individuals of size $n$ should intersect in order to strike a balance between the opportunities to coincide in the common overlap and the amount of unique knowledge each member of the set should hold outside of that area. In other words, there is an intersection/union ratio, $w^{*}$ such that the amount of information shared is maximized for a set of $n$ many individuals. In the Supplementary Information, we show that $w^{*}=1/(n+1)$. In other words, larger sets of individual areas should have a lower intersection/union ratio, according to the convex decreasing curve in dashed lines in Figure 4. The empirical values of this ratio similarly show a convex decreasing trend for all overlapping sets, in many cases qualitatively fitting the theoretical prediction (Figure 4). Overall, smaller sets of up to five individuals vary more widely in the value of $w$ than larger sets. The qualitative fit between the optimal prediction and the observations is particularly good for some seasons (dry seasons in 2014 and 2017 and wet seasons in 2013, 2016 and 2017), but less so in others (dry season in 2016, wet season in 2015). At fixed $n$, there also appears to be a multimodal distribution of the observed values of $w$, with the lowest values resembling the theoretical prediction and others being higher than predicted, suggesting other processes that are not captured by our theoretical model. In sum, the prediction of a balance between redundant and unique areas is overall consistent with the observed patterns, but overlaps tend to be more redundant than predicted by our model alone.
@@
Figure 4: Scaling relationship between the number of individual core ranges and its intersection/union ratio, $w$. Each dot corresponds to a particular set of a given number of individuals (see Methods for details). The dotted line corresponds to the scaling relationship $w^{*}_{n}=1/(n+1)$. Top row: dry seasons; bottom row: wet seasons.
@@
The previous results on the relevance of $w^{*}$ led us to use it as a reference for constructing simplicial complexes depending on how far a given set of individual core ranges was to the predicted $w^{*}$. Each simplex has an $\alpha$ value calculated from their value of $w$, taking into account how the theoretical optimal prediction $w^{*}$ varies with simplex size, $n$. At low values of $\alpha$, simplicial complexes are constructed only containing sets of individual areas that overlap considerably (i.e. have relatively more redundant spatial interactions), while high values of $\alpha$ implies the additional inclusion of sets of individuals that overlap less (i.e. have relatively more unique spatial interactions).
@@
When we use $\alpha$ as a filtration parameter, including in the simplicial complex only those simplices that are below the $\alpha$ threshold, we find evidence of structural holes at various dimensions (Figure 5). Multiple connected components, representing holes in the lowest dimension, persist for values higher than $\alpha=2$, especially in the dry seasons. In some circumstances, only when including the overlaps with the lowest level of redundancy do the simplicial complexes become fully connected. The dry season of 2017, in particular, shows 3 connected components for all values of $\alpha$. Bi-dimensional holes ($\beta_{1}$) and voids ($\beta_{2}$) are common when including interactions beyond $\alpha=3$ and often persist for several units of $\alpha$. Holes of dimension 4 ($\beta_{3}$) only appear in the dry seasons of 2015 and 2016, while two long-persisting holes of dimension 5 ($\beta_{4}$) appear in the wet season of 2012. In the dry season of 2015, holes of four different dimensions ($\beta_{0}$ to $\beta_{3}$) appear at some $\alpha$ values, representing a particularly rich spatial structure across the filtration. Overall, these results provide compelling evidence that there are many areas occupied uniquely by subsets of individuals of various sizes, and, if we interpret these overlaps as indications of information sharing, there is potentially a large amount of complementary information being shared. Pockets of unique knowledge exist among some sets of individuals even as we include the corresponding spatial overlaps with less redundant information (higher $\alpha$).
@@
Figure 5: Persistence barcodes for the simplicial complexes in each season, including the simplices of individual core ranges formed by decreasingly redundant interactions, as $\alpha$ increases. Top row: dry seasons; bottom row: wet seasons. Each barcode shows, for increasing values of $\alpha$, the persistence of holes of different dimension as specified in the legend, corresponding to the different Betti numbers ($\beta$). The beginning and end of each bar correspond to the values of $\alpha$ at which the hole appears and disappears, respectively. By definition, at $\alpha=0$, each individual in the sample is a connected component on its own. While one connected component will persist for all values of $\alpha$, others which are initially present become joined with others and disappear, as $\alpha$ increases and less redundant interactions are used to join subsets of individual core ranges into simplices (see also Figure 1 for an explanation of the filtration and the resulting persistence barcodes).
@@
A qualitative exploration of the persistence barcodes in Figure 5 suggests that the dry season contained more topological holes, which were also more persistent, providing stronger evidence for knowledge complementarity in the dry seasons. To quantify the patterns in each persistence barcode, we developed a filtration complementarity index, which measures the interval over $\alpha$ for which different holes last over the filtration, and gives more weight to those of higher dimension. The values of this index show no clear evidence for simplicial complexes from the dry seasons to have higher complementarity than for those of the wet seasons (ANOVA $F_{1}$=2.2, $p=$0.17; Figure SI4), although our power to detect an effect is weak given our small sample size of 6 years. The pattern is similar for the correlation with the variation in fruit abundance across fortnights for each season. While seasons with higher variation in fruit abundance have slightly larger values of complementarity on average (Figure SI1), we did not find statistical support for this pattern ($\rho$ = 0.42 and $p$=0.16).
@@
To examine the structure of the simplicial complexes, we utilise the maximal simplicial degree centrality, a generalisation of degree centrality in dyadic networks. Across the majority of seasons, we found a clear sigmoidal relationship between the centrality of a simplex and its size. Maximal degree centrality increases gradually with size of the simplices and then more steeply, levelling off or decreasing after the simplex is of size 5. This implies that simplices of this size are involved in a large and variable number of spatial interactions with other simplices, while above a certain size they do not seem to increase their involvement in more spatial interactions with other simplices. We found no clear trends in the relationship between the simplicial degree centrality of a simplex and the average age, the proportion of males or the proportion of immigrant individuals in the individuals comprising it (see Supplementary Information, Figures SI2-4).
Figure 6: Maximal simplicial degree centrality values for simplices of different size. Each dot corresponds to a simplex formed by the intersection of individual core ranges assuming $\alpha\leq 4$. Dashed black line corresponds to a continuous sigmoid function fitted to the centrality values using the non-linear least squares method from the Scipy package in Python [SciPy2020]. Top row: dry seasons; bottom row: wet seasons.

### v3

# Uncovering complementary information sharing in spider monkey collective foraging using higher-order spatial networks

@@ Abstract
Collectives are often able to process information in a distributed fashion, surpassing each individual member’s processing capacity. In fission-fusion dynamics, where group members come together and split from others often, sharing complementary information about uniquely known foraging areas could allow a group to track a heterogenous foraging environment better than any group member on its own. We analyse the partial overlaps between individual spider monkey core ranges, which we assume represent the knowledge of an individual during a given season. Sets of individuals with complementary overlaps are identified, showing a balance between redundantly and uniquely known portions, and we use simplicial complexes to represent these higher-order interactions. The structure of the simplicial complexes shows holes in various dimensions, revealing complementarity in the foraging information that is being shared. We propose that the complex spatial networks arising from fission-fusion dynamics allow for adaptive, collective processing of foraging information in dynamic environments.
@@ Introduction
A driver of the structure and dynamics of networked systems is their adaptive capacity to collectively process information in ways that none of the components would be able to alone [1]. A prevalent feature of this information processing is its distributed nature: different components complement one another in the information they contain, which they aggregate via the network of interactions. This complementarity in information aggregation is a hallmark of collectively intelligent systems [2, 3].
@@
In social animals, the sharing of information about resource availability between group members has long been considered one of the benefits of living in groups [4, 5, 6]. Information transfer between individuals allows the group as a whole to gain better knowledge of the location of available food sources than any individual would gain on its own, especially in heterogenous, dynamic environments [4, 7]. This information transfer need not be through the use of active signals, and may simply require that naïve individuals follow knowledgeable ones to find food sources that the latter know uniquely (e.g. [8, 9, 10]). In species where group members disperse over wide foraging areas, knowledge could be distributed among all of them, such that each may act as naïve or knowledgeable with respect to different food sources [8, 11]. Sharing foraging information thus may allow the benefits of being in a group to outweigh the costs of competition caused by foraging with others [4].
@@
Fission-fusion dynamics are a property of many animal groups [12, 13] that occupy a common area but are often divided into subgroups that frequently come together and split from each other. In species with a high degree of these dynamics, the entire group is never found together, and subgroups of varying sizes spread over large areas, travelling independently of other subgroups [14] but also coming together often [15]. In terms of group-decision making, fission-fusion dynamics can be seen as the result of a lack of consensus between individuals who may have different priorities on their activities [16]. However, in heterogenous, dynamic environments, this lack of consensus may have advantages in terms of distributed foraging and information gathering: individuals may forage in different regions of the group home range, while in subgroups with different combinations of individuals [17]. Thus, a subgroup will often contain individuals who have been foraging in (and perhaps have knowledge about) different regions of their home range.
@@
Spider monkeys (Ateles spp) are frugivorous primates living in tropical forests, where fruit can be found in ephemeral patches that vary in time and space depending on the tree species [18]. They have a high degree of fission-fusion dynamics [19] in terms of the variation in subgroup size [15] and composition [17], and they occupy partially overlapping home ranges [20]. In previous work [11], we showed that spider monkeys share information about newly discovered fruiting trees. Focusing our observations on relevant trees during their whole fruiting period, we showed that individuals that were naïve about the presence of fruit in the focal tree tended to arrive with other group members that already knew about it. Using a random arrival null model, we also showed that the group as a whole finds out about the fruiting tree in fewer visits than if each individual had to find it on its own [11].
@@
Partly inspired by those results, Falcón-Cortés et al. [21] used an agent-based model to explore the effects of information sharing and memory of visited patches on the collective estimation of food patch quality. When individual agents remember previously visited patches, move toward them depending on their quality, and copy others’ visiting patterns if not finding food on their own, the group as a whole is able to compute the relative rank of food patches in the environment, concentrating its activity in proportion to the size of patches and focusing foraging efforts around the largest patches [21]. These studies show that sharing complementary information about heterogenous environments can effectively lead to an improvement in the group’s overall knowledge, and thus could be considered a form of collective information processing, sensu [22].
@@
Here we assume that information sharing takes place between individuals, in the context of a fluid grouping pattern where knowledgeable and naïve individuals coincide temporarily in the same subgroups and locations, being able to share information with others about areas they uniquely know. We also assume that an individual’s core range represents, for a given season, the area where it knows the location of available fruiting trees relatively well. This assumption is supported by the species’ wide daily travel distance (between 500 and 4500m with means between 1750 and 3311m in different study sites [23]) when an individual visits around 10 different fruiting trees [24]. Individual core ranges vary across seasons as would be predicted based on fruit availability, with dry seasons showing larger and less overlapping core ranges than wet seasons [20].
@@
The way in which individual core ranges overlap shows that, normally, at least two individuals (but not the whole group) coincide in most portions of a group’s range [20], such that range overlaps can be thought of as higher-order interactions between two or more individuals. Therefore, typical network approaches to describe the full space-sharing structure of the group are inadequate due to their inability to capture multibody (non-dyadic) interactions between individuals [25, 26]. Simplicial sets are increasingly applied to model such scenarios where non-dyadic interactions are key features of network structure [27, 28, 29, 30]. Simplicial sets are generalisations of ordinary networks which allow for arbitrarily many nodes or individuals to be connected by a single higher-order edge called a simplex (Figure 1). For spatial problems, a particularly relevant class of simplicial structure is the simplicial complex [25] which imposes the additional assumption of downward closure; meaning that the simplicial set must contain all possible nested lower-order simplices of any simplex [26, 31, 32]. For example, a simplicial complex that includes a triadic connection between three individuals must also contain the three dyadic connections between these individuals. This assumption of downward closure provides a suite of additional computational tools for studying higher-order structure [32].
@@
Simplicial complex structure is often described through the computation of Betti numbers. There is a Betti number for each dimension, denoted $\beta_{i}$, which gives the number of holes of a particular simplicial complex in the $i$-th dimension. For instance, $\beta_{0}$ is the number of connected components, $\beta_{1}$ the number of loops and $\beta_{2}$ the number of voids [32]. For higher dimensions, we cannot visualise the holes, so the Betti numbers are especially useful to describe higher dimensional topological structure. In the context of a higher-order network structure of spatial overlaps, the topological holes described by the Betti numbers might result from individuals, or subsets of them, using areas uniquely with respect to others while maintaining some degree of redundancy in the common intersection. Thus, Betti numbers may characterise complementarity in the information possessed by different subsets (represented here by simplices). In addition, simplicial centrality measures [31] allow the identification of simplices that are well-connected to others in the simplicial complex.
@@
We hypothesise that, by pooling complementary information about the location of feeding trees, the group as a whole acquires a more complete knowledge of a complex, dynamic environment during a given season than any individual on its own. In particular, given our knowledge of spider monkey socio-spatial ecology, we expect to find evidence of complementarity in the structure of the simplicial complex representing the overlaps of the individual core ranges. Our strategy, outlined in Figure 1, involves first calculating individual core ranges and their relative intersections. We assume that complementary foraging information can be captured through the intersection/union ratio for any given set of individual core areas, representing a balance between redundant and uniquely known information. The intersection is where the set can coincide to share information about those areas known uniquely by subsets of individuals in the rest of their union.
@@
We use varying thresholds of this intersection/union ratio to connect sets of individuals with decreasingly redundant overlaps into simplicial complexes. Then, we explore the structure of these complexes, searching for holes in various dimensions (given by the Betti numbers) and their persistence for different thresholds of redundancy. We interpret these holes and their persistence as evidence of complementarity in the foraging information sharing by the whole network. We also compare the simplicial complexes between the dry and wet seasons, as the uncertainty about the foraging environment may be greater in the former. We expected to find evidence for more complementarity in the information being shared during the dry seasons. Finally, we analyse the centrality patterns in the structure of the simplicial complex, to identify those sets of individuals that are key in terms of sharing foraging information with others by participating in more simplices.
@@
Figure 1: Strategy to analyse spatial overlaps in individual core ranges as higher-order interactions in a simplicial complex. (a) Four hypothetical areas from individuals AA, $B$, $C$ and $D$ overlap to different extents. The balance of a spatial interaction between a set of any size, $n$, is defined by the intersection/union ratio $w$, which in our idealised model, scales with $n$ as $w^{*}=1/(n+1)$ (see Methods for details). We use a filtration index, $\alpha$, which describes how far from $w^{*}$ a given $w$ is, to filter the sets of individual core ranges, from those with highly redundant spatial interactions (low $\alpha$) to those with less redundant ones (higher $\alpha$), as outlined in detail in the Methods. In this example, the areas AA, $B$ and $C$ have a more redundant spatial interaction than the areas AA, $B$ and $D$ due to their higher intersection/union ratio. Thus, the 2-simplex $[A,B,C]$ is created at lower $\alpha$ values than the 2-simplex $[A,B,D]$. (b) Different configurations of the resulting simplicial complex as $\alpha$ increases (going from left to right, adding links depending on the redundancy of the spatial interaction). Assuming no interactions would represent these four areas as nodes in an unconnected graph. As we incorporate spatial interactions according to a decreasing intersection/union ratio, we introduce links that join the initially disconnected nodes. First the link between $A$ and $B$, who intersect the most, then $B$ and $C$, later AA and $C$. As these simplices appear and close a circuit, they leave a hole of dimension 1 which is filled by a 2-simplex when our filtering includes the interaction between the three of them. Finally, when incorporating the smaller and less redundant spatial interaction between $D$ and the rest, we obtain a 3-dimensional simplicial complex which contains all possible simplices within it, including the three-dimensional space inside the tetrahedron, representing the four-way interaction between all nodes. Above each simplicial complex we note the simplices it contains (in brackets), and the order of the resulting simplicial complex, given by the highest-order simplex it contains. Below each simplicial complex are its Betti numbers $\beta_{0}$ and $\beta_{1}$. (c) The corresponding barcode diagram representing the persistence of the four connected components (red bars) as we incorporate less redundant and more unique interactions and the persistence of a 1-dimensional hole (blue bar). The Betti numbers at each step in the filtration correspond to the number of red ($\beta_{0}$) and blue ($\beta_{1}$) bars present from that step onward along the filtration.
@@ Results
Figure 2: Maps of overlaid individual core ranges, for different years and seasons (top row: dry seasons; bottom row: wet seasons). Each polygon filled with a different colour shade corresponds to an individual (scale bar in the lower right). Below each map is the number of individual core ranges overlaid, as well as the area comprised by their union, in hectares. The northwestern part of the group’s home range borders the Punta Laguna lake (see [20] for more information about the spatial ecology of spider monkeys at the study site).
Figure 3 shows the density distribution of individual core ranges for the different years and seasons. Individual ranges differed significantly between years and seasons (ANOVA interaction between years and seasons, $F_{1,135}$=1.94, pp=0.09; year: pp<0.001; season: $p$ <0.001). During a given year, individual spider monkeys occupied larger core ranges during the dry season than during the wet season. This is particularly evident in 2014, 2015 and 2016.
@@
We developed a theoretical prediction of the extent to which the core ranges of a set of individuals of size $n$ should intersect in order to strike a balance between the opportunities to coincide in the common overlap and the amount of unique knowledge each member of the set should hold outside of that area. In order to develop this prediction (detailed in the Supplementary Information), we first assume that a particular spatial overlap structure in a given subset of individual core ranges has a degree of complementarity depending on the amount of information that can be transferred between individuals. This information transfer, in turn, is proportional to the area known uniquely by some individuals but not others, and to the opportunities they have to coincide in a common area in order to share their uniquely known information. The higher the transfer of information, the more complementary their relative use of areas is. We also assume that individuals are distributed uniformly within their own areas, and independently from each other. Then we constructed an objective function describing this transfer of information and optimised it under the additional assumption of an homogeneous foraging ability in the population, i.e., individual core ranges of identical sizes (see Supplementary Information for more details). We found that the optimal intersection/union ratio for a set of $n$-many individuals is $w^{*}=1/(n+1)$.
The above relationship predicts that larger sets of individual areas should have a lower intersection/union ratio, according to the convex decreasing curve shown in Figure 4. The empirical values of this ratio similarly show a convex decreasing trend for all overlapping sets, in many cases qualitatively fitting the theoretical prediction (Figure 4). Overall, smaller sets of up to five individuals vary more widely in the value of $w$ than larger sets. The qualitative fit between the optimal prediction and the observations is particularly good for some seasons (dry seasons in 2014 and 2017 and wet seasons in 2013, 2016 and 2017), but less so in others (dry season in 2016, wet season in 2015). At fixed $n$, there also appears to be a multimodal distribution of the observed values of $w$, with the lowest values resembling the theoretical prediction and others being higher than predicted, suggesting other processes that are not captured by our theoretical model. In sum, the prediction of an optimal value of ww*, representing a complementary balance between redundant and unique information, is overall consistent with the observed patterns, but overlaps tend to be larger (i.e. more redundant) than predicted by our model alone.
@@
Figure 4: Scaling relationship between the number of individual core ranges and its intersection/union ratio, $w$. Each dot corresponds to a particular set of a given number of individuals (see Methods for details). The solid black line corresponds to the scaling relationship $w^{*}_{n}=1/(n+1)$. Top row: dry seasons; bottom row: wet seasons.
@@
The previous results on the relevance of $w^{*}$ led us to use it as a reference for constructing simplicial complexes depending on how far a given set of individual core ranges was to the predicted $w^{*}$. Each simplicial complex has an $\alpha$ value calculated from their value of $w$, taking into account how the theoretical optimal prediction $w^{*}$ varies with number of individuals, $n$. At low values of $\alpha$, simplicial complexes are constructed only containing sets of individual areas that overlap considerably (i.e. have highly redundant spatial interactions), while high values of $\alpha$ imply the additional inclusion of sets of individuals that overlap less (i.e. have relatively more unique spatial interactions).
@@
When we use $\alpha$ as a filtration parameter, including in the simplicial complex only those simplices that are below the $\alpha$ threshold, we find evidence of structural holes at various dimensions (Figure 5). Multiple connected components, representing holes in the lowest dimension, persist for values higher than $\alpha=2$, especially in the dry seasons. In some circumstances, only when including the overlaps with the lowest level of redundancy do the simplicial complexes become fully connected. The dry season of 2017, in particular, shows 3 connected components for all values of $\alpha$. Bi-dimensional holes ($\beta_{1}$) and voids ($\beta_{2}$) are common when including interactions beyond $\alpha=3$ and often persist for several units of $\alpha$. Holes of dimension 4 ($\beta_{3}$) only appear in the dry seasons of 2015 and 2016, while two long-persisting holes of dimension 5 ($\beta_{4}$) appear in the wet season of 2012. In the dry season of 2015, holes of four different dimensions ($\beta_{0}$ to $\beta_{3}$) appear at some $\alpha$ values, representing a particularly rich spatial structure across the filtration.
If we assume that a hole of any dimension in the simplicial complex is the result of unique areas being used by some individuals in the group but not by others, then we can interpret holes and their persistence over different values of the filtration parameter as evidence of complementarity in the foraging information shared by these subsets. The existence of these holes implies that pockets of unique knowledge may exist among some sets of individuals even as we include spatial overlaps with less redundant information (higher $\alpha$). Overall, these results provide compelling evidence that there are many areas occupied uniquely by subsets of individuals of various sizes, and, if we interpret these patterns of overlaps as indications of partial information sharing, there is potentially a large amount of complementary information being shared.
@@
Figure 5: Persistence barcodes for the simplicial complexes in each season, including the simplices of individual core ranges formed by decreasingly redundant interactions, as $\alpha$ increases. Top row: dry seasons; bottom row: wet seasons. Each barcode shows, for increasing values of $\alpha$, the persistence of holes of different dimension as specified in the legend. The number of holes of each dimension $i$ at a given filtration step corresponds to each Betti number ($\beta_{i}$). The beginning and end of each bar correspond to the values of $\alpha$ at which the hole appears and disappears, respectively. By definition, at $\alpha=0$, each individual in the sample is a connected component on its own. While one connected component will persist for all values of $\alpha$, others which are initially present become joined with others and disappear as $\alpha$ increases and less redundant interactions are used to join subsets of individual core ranges into simplices (see also Figure 1 for an explanation of the filtration and the resulting persistence barcodes). Note that our persistence barcodes do not finish with trivial structures (containing each possible simplex) due to the inclusion of a maximal simplex size, $n_{\text{max}}$.
@@
A qualitative exploration of the persistence barcodes in Figure 5 suggests that the dry season contained more topological holes, which were also more persistent, providing stronger evidence for knowledge complementarity in the dry seasons. To quantify the patterns in each persistence barcode, we developed a filtration complementarity index, which measures the interval over $\alpha$ for which different holes last over the filtration, and gives more weight to those of higher dimension. The values of this index show no clear evidence for simplicial complexes from the dry seasons to have higher complementarity than for those of the wet seasons (ANOVA $F_{1}$=2.2, p=p=0.17; Figure SI4), although our power to detect an effect is weak given our small sample size of 6 years. The pattern is similar for the correlation with the variation in fruit abundance across fortnights for each season. While seasons with higher variation in fruit abundance have slightly larger values of complementarity on average (Figure SI1), we did not find statistical support for this pattern ($\rho$ = 0.42 and pp=0.16).
@@
To examine the structure of the simplicial complexes, we utilise the maximal simplicial degree centrality, a generalisation of degree centrality in dyadic networks [31]. Across the majority of seasons, we found a clear sigmoidal relationship between the centrality of a simplicial complex and its size. Maximal degree centrality increases gradually with size of the simplices and then more steeply, levelling off or decreasing after the simplex is of size 5. This generally implies that simplices containing more individuals are involved in a larger and more variable number of spatial interactions with other simplices, while above a certain size there is a plateau in this effect. We found no clear trends in the relationship between the simplicial degree centrality of a simplex and the average age, the proportion of males or the proportion of immigrant individuals in the individuals comprising it (see Supplementary Information, Figures SI2-4). Note that the increasing trend of centrality with simplex size is consistent when varying the parameter $n_{\text{max}}$ (see Supplementary Information, Figures SI5-9), implying that it is indeed larger simplices which are more central and not those of some intermediate size which we could not observe due to our choice of $n_{\text{max}}=6$.
Figure 6: Maximal simplicial degree centrality values for simplices of different size. Each dot corresponds to a simplex formed by the intersection of individual core ranges assuming $\alpha\leq 4$. The dashed black line corresponds to a continuous sigmoid function fitted to the centrality values using the non-linear least squares method from the Scipy package in Python [33]. Top row: dry seasons; bottom row: wet seasons.